ISS On-Orbit Status 06/10/09
All ISS systems continue to function nominally, except those noted previously or below. Today’s sleep/wake cycle:
- Sleep today – 8:20am; wake – 2:00am tomorrow (back on normal)
The Russian Orlan EVA-23 “internal” spacewalk by CDR Gennady Padalka & FE-1 Michael Barratt began at 2:55am EDT and lasted 12 min, concluding successfully at 3:07am EDT. Attitude control authority was handed over to Russian MCS (Motion Control System) thrusters at 12:00am and returned to U.S. momentum management at ~3:30am. [Connected to Orlan life support umbilicals (i.e., no backpacks), Gennady & Mike worked inside the depressurized SM PkhO (Service Module Transfer Compartment), removing the flat hatch cover of its +Y-axis docking assembly and replaced it with the standard conical docking cone cover. The conical cover is required for berthing of the Russian MRM2 (Mini Research Module 2) scheduled to arrive in November this year on a Proton, for which the previous EVA-22, on 6/5, installed docking antennas, a docking target and electrical connectors on the exterior of the SM. MRM2 will provide extra docking capability for Soyuz vehicles, thus freeing up the DC1 port for Progress cargo carriers. This allows for a more effective ISS roll control using the Progress propulsion system. After the EVA, the MLI (Multilayer Insulation) of the removed flat hatch cover began flaking off and became air-born. Protected with goggles and dust masks, the crew vacuumed up the particles.]
After wakeup last night at ~7:00pm EDT, Mike Barrat & Koichi Wakata continued their new recording rounds for the experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) logging data from their Actiwatch to the HRF-1 (Human Research Facility 1) laptop as part of a week-long session. It is the third for Mike, the fourth for Koichi. [To monitor the crewmember’s sleep/wake patterns and light exposure, Mike & Koichi wear a special Actiwatch device which measures the light levels encountered by them as well as their patterns of sleep and activity throughout the Expedition and use the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days.]
Padalka & Barratt took the standard pre-EVA MO-9 “Urolux” urine biochemistry test before having breakfast.
FE-4 Thirsk worked on the ARED (Advanced Resistive Exercise Device), removing the old pulley rope and replacing it with a new Vectran cord, carefully splicing it through the pulley system with the assistance of a crewmate. [It was a one-hour activity, plus an additional 30-min of assistance.]
Later, Bob Thirsk completed the periodic maintenance & visual inspection of the ARED and its VIS (Vibration Isolation System) guide rails & rollers, greasing the Y- and Z-axis rails & rollers and also evacuating its cylinder flywheels to maintain proper vacuum condition and sensor calibration.
In the U.S. Lab, Bob started (later terminated) another 5-hr automatic sampling run, the seventh, with the new EHS GC/DMS (Environmental Health System Gas Chromatograph/Differential Mobility Spectrometer), also known as AQM (Air Quality Monitor), controlled with “Sionex” expert software from the SSC-4 (Station Support Computer 4) laptop. [The AQM demonstrates COTS (Commercial Off-the-Shelf) technology for identifying volatile organic compounds, similar to the VOA (Volatile Organics Analyzer). Today’s data will again to be compared with VOA and GSC (Grab Sample Container) measurements. This evaluation will continue over the course of several months as it helps to eventually certify the GC/DMS as nominal CHeCS hardware.]
FE-3 Romanenko had 2 hrs set aside for unloading cargo from the Soyuz TMA-15/19S spacecraft at the FGB nadir port.
Later, Romanenko completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]
After conclusion of the EVA-23 at ~3:07am, Gennady & Mike, assisted by Roman –
- Repressurized the SM PkhO transfer compartment,
- Conducted their second MO-9 “Urolux” biochemical urine test,
- Reset STTS communications in the SM/PkhO,
- Re-installed the air duct through the PkhO hatch,
- Restored systems configurations in the SM to pre-EVA conditions, and
- Set up the Orlan-MK suits, umbilicals and BSS interface units for drying out.
Also after the spacewalk, Thirsk & DeWinne in turn conducted their second sessions with the experiment BISE (Bodies in the Space Environment), complete with videocam coverage, investigating the relative contributions of internal and external cues to self-orientation during and after zero-G exposure. After setting up the camcorder for recording the activity, configuring the “Neurospat” hardware and activating the BISE software on its A31p SSC (Station Support Computer) laptop, Bob & Frank each then had ~20-25min for completing the experiment protocol. [The CSA (Canadian Space Agency)-sponsored BISE experiment studies how astronauts perceive Up and Down in microgravity. The specific objective of the BISE project is to conduct experiments during long-duration microgravity conditions to better understand how humans first adapt to microgravity and then re-adapt to normal gravity conditions upon return to earth. This experiment involves comparisons of preflight, flight, and post-flight perceptions and mental imagery, with special reference to spaceflight-related decreases in the vertical component of percepts. The test involves having subjects view a computer screen through a cylinder that blocks all other visual information. The astronauts are being presented with background images with different orientations relative to their bodies.]
The FE-4 performed the regular controlled shut-down of the EHS VOA (Environmental Health System-Volatile Organic Analyzer), with the ground power-cycling its RPC-3 (Remote Power Controller 3), part of RPCM (RPC Module) LAD42B_A.
Thirsk also conducted the standard sensor calibration and check on the CSA-O2
(Compound Specific Analyzer-Oxygen) units #1043 & #1059, delivered on Flight 1J, then take readings in the Lab.
FE-2 Wakata meanwhile conducted a major task review of the upcoming Robotics/MSS (Mobile Service System) and SSRMS (Space Station Remote Manipulator System)
activities during STS-127/2JA. [OBT (On-board Training) for MSS should be performed not earlier than 10 days prior to task. Robotics skills degrade over time, similar to piloting skills. Also, the complexity and nuances of the MSS system require system knowledge refreshing gained either through system check-outs or OBTs.]
Wakata undertook another periodic relocation of the TEPC (Tissue Equivalent Proportional Counter) detector assembly, the primary radiation measurement tool in the ISS, moving it in the SM to a location forward of the TORU control panel (used last year), then taking documentary photography, downloaded to SSC-13. [Due to somewhat higher-than-expected readings seen in the recent TEPC data, only the TEPC detector was relocated today. Comparisons between data gathered from this new location and data gathered from this same location in January 2008 will allow radiation specialists to determine the validity of these readings.]
Koichi also conducted a run with the MedOps experiment WinSCAT (Spaceflight Cognitive Assessment Tool for Windows), his third onboard session, by logging in on the MEC (Medical Equipment Computer) laptop and performing the psychological evaluation exercise on the PC-based WinSCAT application. [WinSCAT is a monthly time-constrained questionnaire test of cognitive abilities, routinely performed by astronauts aboard the ISS every 30 days before or after the PHS (periodic health status) test or on special CDR's, crewmembers or flight surgeons request. The test uses cognitive subtests that measure sustained concentration, verbal working memory, attention, short-term memory, spatial processing, and math skills. The five cognitive subtests are Coding Memory - Learning, Continuous Processing Task (CPT), Match to Sample, Mathematics, and Coding Delayed Recall. These WinSCAT subtests are the same as those used during NASA’s long-duration bed rest studies.]
Roman completed the daily IMS (Inventory Management System) maintenance, updating/editing its standard “delta file” including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).
In the Kibo JPM (JEM Pressurized Module), FE-4 Thirsk supported JAXA ground-controlled payload work (FACET) by turning off the PLT (Payload Laptop Terminal) which he had activated yesterday.
Koichi prepared another JAXA Astronaut’s Short Report (reflecting “ordinary life” in space, intended to attract children and general public's attention for future human space exploration), followed by another J-ASTRO Educational Report, to show the significance of living in space and to respond to some questions by children and from some Japanese intellectuals. [Both reports were later submitted to NASA CB (Astronaut Office), JAXA and NASA PAO representatives.]
The FE-2 also filled out the regular weekly FFQ (Food Frequency Questionnaire) on the MEC, his 13th
. [On the FFQs, NASA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. The FFQ is performed once a week to estimate nutrient intake from the previous week and to give recommendations to ground specialists that help maintain optimal crew health. Weekly estimation has been verified to be reliable enough that nutrients do not need to be tracked daily.]
Thirsk reduced cooling on the CHeCS (Crew Health Care Systems) rack by demating the ITCS MTL (Internal Thermal Control System Moderate Temperature Loop) return umbilical at the rack’s UIP (Utility Interface Panel, loc. LAB1S4) to increase coolant flow for payload use. [The hose had been mated by Koichi on 6/7 to provide cooling to the rack.]
Frank & Bob had another 2.5 hrs for gathering, prepacking & staging cargo itemized in an uplinked list for return on STS-127/Endeavour (2J/A) later this month.
The crew (except for the spacewalkers) completed their regular daily 2.5-hr. physical workout program on the CEVIS cycle ergometer (FE-4, FE-5), TVIS treadmill (FE-2), ARED advanced resistive exercise device (FE-2, FE-3, FE-4, FE-5) and VELO cycle ergometer with bungee cord load trainer (FE-3). [On the CEVIS, the actual loads remain slightly lower than the commanded loads, but this was expected. A manual correction of the pertinent calibration coefficient via the control panel touch screen will be done at a later time when the new value has been determined.]
Later, Bob transferred the exercise data file to the MEC for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on ARED, followed by their erasure on the HRM storage medium (done six times a week). OGS Status Update:
The U.S. OGS (Oxygen Generator System) in the Lab has been exhibiting a rise in delta-pressure across the pump package assembly. The “shutdown” limit set in the software is 25 psi, and the data are trending exponentially upward toward this value within the last week. Engineers are working on procedures to clean the filter in the system in case 2J/A does not launch on 6/13 (a spare pump and filter are manifested on Endeavour). If launched on time, the plan is to replace the filter during the 2J/A mission and return the old filter. If the filter change-out is unsuccessful in reducing the rising delta-P, then the problems could be in the water ORU (Orbit Replaceable Unit). There is a spare ORU onboard but the R&R is much more intrusive than the pump-and-filter change. As of today, the procedure will be to shut down the OGS when delta-P exceeds 23 psi (1190 mmHg). OGS is now being moded to 100% production for a maximum of 48 hrs, to get extra O2
into the ISS interior before OGS shutdown. After 48 hrs, OGS is returned to 30% production, staying there until pump delta-P exceeds 23 psi. Changing to 100% allows maximum production over the final few days that OGS is operational (prior to the R&R). TVIS Update:
FE-2 Wakata completed the TVIS ACO (Activation & Checkout) this morning and recounted that everything felt and sounded nominal. Koichi reported that there were no anomalous sounds during the shake test on all four stabilizers. Additionally, during the exercise session, the TVIS was stable, unlike previous exercise sessions where oscillations were observed. Data and HD (High Definition) video recorded were downlinked for specialists to review prior to giving the crew a Go to exercise. STS-127/Endeavour (2J/A) Forecast:
Weather forecast has improved since yesterday and looks very favorable for launch: Probability of KSC weather prohibiting launch: 10%.
CEO (Crew Earth Observation) photo targets uplinked for today were Bigach Impact Crater, Kazakhstan (looking slightly left of nadir over this 8 km in diameter impact structure located the west of Lake Zhaysang. The crater is visible on the landscape as a rough ring of disturbed rock surrounding an almost flat interior. Nadir viewing, overlapping frames taken along track were requested),
and Baikonur Cosmodrome, Kazakhstan (light cloud cover could have been present over the Cosmodrome at the time of the ISS overpass. The crew should have been able to see the target just to the right of the orbital track. Recommendations were to take overlapping, nadir-viewing frames as during approaching, passing over, and departing the target area to ensure capturing imagery of the target)
CEO photography can be studied at this “Gateway” website: http://eol.jsc.nasa.gov
(as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone). ISS Orbit (as of this morning, 8:16am EDT [= epoch])
Mean altitude -- 348.6 km
Apogee height – 354.8 km
Perigee height -- 342.4 km
Period -- 91.51 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0009228
Solar Beta Angle -- -28.4 deg (magnitude decreasing)
Orbits per 24-hr. day -- 15.73
Mean altitude loss in the last 24 hours -- 71 m
Revolutions since FGB/Zarya launch (Nov. 98) -- 60490 Significant Events Ahead (all dates Eastern Time, some changes possible!)
06/13/09 -- STS-127/Endeavour/2J/A launch - JEM EF, ELM-ES, ICC-VLD (7:12am)
06/29/09 -- STS-127/Endeavour/2J/A landing (12:18am EDT, KSC)
07/17/09 – Progress M-02M/33P undock & deorbit
07/20/09 -- Soyuz TMA-14/18S relocation (from SM aft to DC1)
07/24/09 -- Progress 34P launch
07/26/09 -- Progress 34P docking (SM aft)
08/07/09 -- STS-128/Discovery/17A – MPLM (P), LMC (~8:49am EDT)
09/01/09 -- H-IIB (JAXA HTV-1) launch – tentative
09/07/09 -- H-IIB (JAXA HTV-1) berth
09/30/09 -- Soyuz TMA-16/20S launch
10/02/09 -- Soyuz TMA-16/20S docking (SM aft, until MRM-2 w/new port)
10/08/09 -- H-IIB (JAXA HTV-1) unberth
10/11/09 – Soyuz TMA-14/18S undock
10/15/09 -- Progress 35P launch
11/10/09 -- 5R/MRM-2 (Russian Mini Research Module 2) on Proton -- tentative
11/12/09 -- STS-129/Atlantis/ULF3 - ELC1, ELC2
12/07/09 -- Soyuz TMA-17/21S launch
12/26/09 -- Progress 36P launch
02/03/10 -- Progress 37P launch
02/04/10 -- STS-130/Endeavour/20A – Node-3 + Cupola
03/18/10 -- STS-131/Discovery/19A – MPLM(P), LMC
04/02/10 -- Soyuz TMA-18/22S launch
04/27/10 -- Progress 38P launch
05/14/10 -- STS-132/Atlantis/ULF4 – ICC-VLD, MRM-1
05/29/10 -- Soyuz TMA-19/23S launch
06/25/10 -- Progress 39P launch
07/29/10 -- STS-133/Endeavour/ULF5 – ELC4, MPLM
08/11/10 -- Progress 40P launch
09/16/10 -- STS-134/Discovery/ULF6 – ELC3, AMS
09/29/10 -- Soyuz TMA-20/24S launch
10/19/10 -- Progress 41P launch
11/??/10 -- ATV2 – Ariane 5 (ESA)
12/??/11 -- 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton